OpenStudy (anonymous):

A 0.4987-g sample of a compound known to contain only carbon, hydrogen, and oxygen was burned in oxygen to yield 0.9267 g of carbon dioxide and 0.1897 g of water. What is the empirical formula of the compound?

OpenStudy (anonymous):

(0.9267 g CO2) / (44.00964 g CO2/mol) x (1 mol C / 1 mol CO2) x (12.01078 g C/mol) = 0.252908 g C (0.1897 g H2O) / (18.01532 g H2O/mol) x (2 mol H / 1 mol H2O) x (1.007947 g H/mol) = 0.0212272 g H Use the Law of Conservation of Mass: (0.4987 g total) - (0.252908 g C) - (0.0212272 g H) = 0.224565 g O Covert to moles: (0.252908 g C) / (12.01078 g C/mol) = 0.0210568 mol C (0.0212272 g H) / (1.007947 g H/mol) = 0.021060 mol H (0.224565 g O) / (15.99943 g O/mol) = 0.0140358 mol O Divide by the smallest number of moles: (0.0210568 mol C) / 0.0140358 mol = 1.500 (0.021060 mol H) / 0.0140358 mol = 1.500 (0.0140358 mol O) / 0.0140358 mol = 1.000 To achieve integer coefficients, multiply by 2, then round to the nearest whole numbers fo find the empirical formula: C3H3O2